# MAPF-HD: Multi-Agent Path Finding in High-Density Environments

### v1

@@ Abstract
Multi-agent path finding (MAPF) involves planning efficient paths for multiple agents to move simultaneously while avoiding collisions. In typical warehouse environments, agents are often sparsely distributed along aisles. However, increasing the agent density can improve space efficiency. When the agent density is high, we must optimize the paths not only for goal-assigned agents but also for those obstructing them. This study proposes a novel MAPF framework for high-density environments (MAPF-HD). Several studies have explored MAPF in similar settings using integer linear programming (ILP). However, ILP-based methods require substantial computation time to optimize all agent paths simultaneously. Even in small grid-based environments with fewer than $100$ cells, these computations can incur tens to hundreds of seconds. These high computational costs render these methods impractical for large-scale applications such as automated warehouses and valet parking. To address these limitations, we introduce the phased null-agent swapping (PHANS) method. PHANS employs a heuristic approach to incrementally swap positions between agents and empty vertices. This method solves the MAPF-HD problem within seconds to tens of seconds, even in large environments containing more than $700$ cells. The proposed method can potentially improve efficiency in various real-world applications such as warehouse logistics, traffic management, or crowd control. Code is available at https://github.com/ToyotaCRDL/MAPF-in-High-Density-Envs.
@@ I Introduction
Industry is increasingly adopting automated guided vehicles (AGVs) and conveyors to enhance efficiency and reduce reliance on human labor. In environments with agents, such as AGVs or pallets on conveyors, efficient collision-free path planning is essential. These path-planning problems for multiple agents are formalized as multi-agent path finding (MAPF) problems [1]. MAPF has been applied to various settings including automated warehouses [2], airport taxiway management [3], automated parking systems [4], and video games [5].
In typical MAPF settings, agents are often sparsely distributed along aisles. However, increasing the agent density can improve space efficiency. For instance, high-density automated warehouses store significantly more goods than lower-density ones, reducing required floor space and operational costs (Fig. 1). Similarly, high-density automated valet parking systems allow vehicles to be parked and retrieved in a smaller area by minimizing the gaps between vehicles.
Despite these space-saving benefits, planning efficient paths using conventional techniques remains challenging. Typically, in high-density environments, only specific agents have active goals at any given time (e.g., moving some items for shipment). The remaining agents, although not assigned to any particular destination, still occupy space and obstruct the movement of the target agents. For the target agents to reach their objectives, the obstructing agents must be relocated efficiently. One approach is to use integer linear programming (ILP) to plan all agent paths simultaneously [6, 7]. However, this approach generally incurs significant computation time because all agent paths are optimized simultaneously by ILP. Another approach is to set temporary goals for obstructing agents and then apply existing MAPF methods. However, this often extends the paths of the target agents because the relocation of the obstructing agents is not optimized. Table I highlights these limitations and the need for a tailored solution to accommodate obstructing agents efficiently.
Figure 1: MAPF in a high-density warehouse environment. Only a subset of target agents (those with matching flags) are assigned goals, whereas the rest act as movable obstacles. In addition to planning paths for target agents to their goals, efficiently relocating obstructing agents is crucial for successful navigation under high-density conditions.
@@
agents agents efficiency
@@
This study refined MAPF for high-density environments (MAPF-HD) and developed an efficient algorithm tailored to this context. This algorithm employs phased heuristic techniques to relocate agents by swapping their positions with empty vertices (null agents). This approach, termed phased null-agent swapping (PHANS), solves MAPF-HD within a few seconds even in a densely packed environment, as shown in Fig. 2.
@@ II-A Optimal Solvers
MAPF involves determining optimal collision-free paths for multiple agents [1, 8]. Conflict-based search (CBS) [9], a widely used optimal solver, plans agents’ paths to their goals (low-level search) and resolves collisions by branching (high-level search). However, CBS faces scalability challenges because MAPF is NP-hard [10]. Researchers have introduced improvements, such as improved CBS, to prune the search more effectively [11]. Other optimal frameworks recast MAPF as a constraint optimization problem. For example, Lam et al. [12] applied the branch-and-cut-and-price ILP method. Similarly, Gómez et al. [13] used answer set programming.
@@ II-B Suboptimal Solvers
Several bounded suboptimal and heuristic algorithms have been developed to improve the scalability of MAPF [14]. Enhanced CBS (ECBS) [15] is a bounded suboptimal solver that employs focal search. To enhance scalability, explicit estimation CBS (EECBS) [16] replaces the focal search in ECBS with explicit estimation search. Priority inheritance with backtracking (PIBT) [17], a decentralized MAPF algorithm, assigns dynamic movement priorities to agents and allows backtracking to resolve conflicts. Unlike static priority planning, PIBT enables agents to inherit and adjust priorities dynamically, preventing deadlocks and improving coordination in dense environments. Backtracking allows agents to reconsider past actions when conflicts arise, thereby ensuring flexible and adaptive navigation.
@@ II-C MAPF in High-Density Environments with Obstructing Agents
Even in dense environments, relocation problems can be resolved when all agents have assigned destinations. However, to move target agents efficiently, relocation of obstructing agents must be optimized.
Okoso et al. [6] introduced the cooperative automated valet parking problem (CoAVP). In CoAVP, agents are classified into three types: entering, departing, and relocating vehicles. Departing vehicles are assigned destinations. Entering and relocating vehicles can be anywhere in the parking lot, and their destinations are also optimized. Their ILP-based method extends the planar graph along the time axis and optimizes the path within the time-expanded network. However, this approach presents two major challenges: First, ILP has high computational complexity and requires hundreds of seconds even in small environments (for example, $15\times 15$ grids). Consequently, this method is impractical in real-world applications. Second, collision types in high-density environments have not been sufficiently considered. To save parking space, agents must be densely positioned. In such environments, two adjacent agents must move in almost perfect synchrony, which may not be realistic. Following conflict [1] should be considered for high-density environments.
@@
This study redefined MAPF in high-density environments and developed a heuristic approach to mitigate the high time complexity associated with ILP.
@@ III Multi-Agent Path Finding Problem in High-Density Environments
This section defines MAPF-HD. Unlike standard MAPF, which focuses on the paths of the target (goal-assigned) agents toward designated goals, MAPF-HD operates in environments with densely located obstructing agents. The objective function of MAPF-HD is to minimize the time steps required for all target agents to reach their goals without collisions. Fig. 3 shows an example of MAPF-HD.
@@ III-A Environment
where $x$ and $x^{\prime}$ are row indices and $y$ and $y^{\prime}$ are column indices. These correspond to the grid environment coordinates. The Manhattan distance $dst(v_{x,y},v^{\prime}_{x^{\prime},y^{\prime}})$ between two points is calculated as follows:
$dst(v_{x,y},v^{\prime}_{x^{\prime},y^{\prime}})=|x-x^{\prime}|+|y-y^{\prime}|.$ (3)
@@ III-B Agent
  • Target agents: $A^{\mathrm{tgt}}={a^{\mathrm{tgt}}_{1},\ldots,a^{\mathrm{tgt}}_{m}}$, with $m>0$. Each target agent $a^{\mathrm{tgt}}_{i}$ is specified by a tuple $\langle s^{\mathrm{tgt}}_{i},g^{\mathrm{tgt}}_{i}\rangle$, where $s^{\mathrm{tgt}}_{i}\in V$ is the start vertex, and $g^{\mathrm{tgt}}_{i}\in V$ is the goal vertex.
  • Obstructing agents: $A^{\mathrm{obs}}={a^{\mathrm{obs}}_{1},\ldots,a^{\mathrm{obs}}_{n}}$, with $n\geq 0$. Each obstructing agent $a^{\mathrm{obs}}_{i}$ is characterized solely by its start vertex, denoted as $\langle s^{\mathrm{obs}}_{i}\rangle$, as it does not have a goal.
@@ III-C Agent Action
Figure 3: Example of the MAPF-HD problem.
The agents avoid collisions with other agents at each time step by managing the two conflicts:
@@
Some studies [9, 16, 17] are based on edge conflict [1] rather than following conflict. If following conflicts are permitted, the agents must be spaced apart, and adjacent agents should move synchronously. However, in agent-dense environments, spacing the agents sufficiently apart may not be feasible. Moreover, perfect synchronization is not always realistic in real environments. Therefore, we adopted following conflicts. Note that following conflicts contain edge conflicts [1].
@@ III-D Solution and Objective
A solution to MAPF-HD is a set of collision-free paths $\{\pi_{1},\ldots,\pi_{m+n}\}$ ensuring that all target agents reach their goals at a certain time step $T$. None of the obstructing agents have set goals, and they are freely positioned. Formally, the problem assigns a path $\pi_{i}=[\pi_{i}[0],\pi_{i}[1],\ldots,\pi_{i}[T]]$ to each agent, which must satisfy the following:
@@
The objective is to minimize makespan $T$, which is defined as the total number of time steps required for all the target agents to reach their goals. In addition to makespan, classical MAPF often use sum-of-costs [1], which is the sum of time steps required by each agent to reach its goal. In real-world applications such as warehouses and valet parking, the emphasis is on the speed of the target agents in reaching their goals, rather than on the total movement effort of all the agents. Makespan captures this operational bottleneck, whereas sum-of-costs would be distorted by the large number of obstructing agents that do not have goals.
@@ IV Proposed Method
Current methods [6, 7] use ILP to optimize the paths of all agents simultaneously, resulting in high computational complexity. To reduce the computational costs, we propose a two-stage heuristic approach termed PHANS. For the target agents, the first stage plans the paths from their starting positions to their goals using the A* algorithm. The second stage sequentially evacuates the obstructing agents from these paths to clear the way for target agents. Section IV-A describes the methods used in both stages through an example with a single target agent, followed by a generalized algorithm for multiple target agents in Section IV-B.
@@ IV-A1 First Stage
To satisfy the second condition, the A* algorithm is modified by introducing an additional heuristic value, $h_{\mathrm{add}}$. This modification incorporates the cost required to wait for obstructing agents to clear the path. The cost function $f$ is computed as the sum of $g,h$, and $h_{\mathrm{add}}$ as follows:
@@
$h_{\mathrm{add}}=\mathrm{max}(0,1+h_{\mathrm{evac}}-g).$ (7)
Here, $h_{\mathrm{evac}}$ estimates the cost of evacuating an obstructing agent by calculating its distance to the nearest empty vertex using (3). If the estimate is accurate and the obstructing agents are evacuated by the time the target agent arrives, $h_{\mathrm{add}}$ is 0. Otherwise, the waiting time is incorporated into $h_{\mathrm{add}}$. These computations are based on the initial positions; however, the empty vertices may move through the evacuation of the obstructing agents.
The proposed heuristic $h+h_{\mathrm{add}}$ is neither admissible nor consistent in general. When $h_{\mathrm{add}}=0$, it reduces to the standard admissible distance estimate. In other cases, it may overestimate the true evacuation cost, and its value can change nonmonotonically between neighboring states. Thus, the heuristic is introduced as a practical cost estimator that reflects the additional effort of relocating obstructing agents in high-density environments, rather than as a guarantee of optimality.
@@ IV-A2 Second Stage
    Plan the path from the null vertex to the obstructing agent (null-agent path), ignoring potential collisions with the other obstructing agents.
@@ IV-B Multiple Target Agents
For multiple target agents, the first stage can be executed simultaneously for all targets. Additionally, the second stage allows the simultaneous evacuation of multiple obstructing agents to enhance efficiency. The pseudocode for this generalized approach is provided in Algorithm 1. We perform sequential A* path planning using prioritized planning (PP) to prioritize moving the target agents farthest from the goal.
Lines 1–4 in Algorithm 1 calculate the distance from the start to the goal for each target agent in $A^{\mathrm{tgt}}$ using (3) and then prioritize the planning of target paths with the longest distances. As resolving conflicts during simultaneous path planning increases computational costs, paths are planned sequentially using PP.
@@
Lines 12–16 assign a null agent to each obstructing agent and plan the path between them. The null-agent assignment proceeds in order from the head of $B$, stopping if no null agents remain. In general, the nearest null agent is assigned to each obstructing agent. However, because target agents are prioritized over obstructing agents, any null agents lying on the target agent’s path between the target agent and the obstructing agent are reserved for the target agent. Therefore, these reserved null agents are not selected as obstructing agents. Lines 18–21 move the target agent after the obstructing agent is evacuated. If there are no collisions, lines 22–25 execute the sequential relocation of the obstructing agents along the paths planned with their assigned null agents.
@@ IV-C How Fast Is PHANS?
Conversely, the methods proposed by Okoso et al. [6] and Makino et al. [7] extended the graph $G=(V,E)$ along the time axis and solved the minimum-cost flow problem through edges for path planning. With a time-axis size of $T$, their worst-case time complexity is $\mathcal{O}\left(2^{T|E|}\right)$.
@@
20:     bagt.path == null agent’s path from bagt.goal to bagt.pos

### v2

@@ Abstract
Multi-agent path finding (MAPF) involves planning efficient paths for multiple agents to move simultaneously while avoiding collisions. In typical warehouse environments, agents are often sparsely distributed along aisles; however, increasing the agent density can improve space efficiency. When the agent density is high, it becomes necessary to optimize the paths not only for goal‑assigned agents but also for those obstructing them. This study proposes a novel MAPF framework for high-density environments (MAPF-HD). Several studies have explored MAPF in similar settings using integer linear programming (ILP). However, ILP-based methods require substantial computation time to optimize all agent paths simultaneously. Even in small grid-based environments with fewer than $100$ cells, these computations can take tens to hundreds of seconds. Such high computational costs render these methods impractical for large-scale applications such as automated warehouses and valet parking. To address these limitations, we introduce the phased null-agent swapping (PHANS) method. PHANS employs a heuristic approach to incrementally swap positions between agents and empty vertices. This method solves the MAPF-HD problem within a few seconds, even in large environments containing more than $700$ cells. The proposed method has the potential to improve efficiency in various real-world applications such as warehouse logistics, traffic management, and crowd control.
@@ I Introduction
Industry is increasingly adopting automated guided vehicles (AGVs) and conveyors to enhance efficiency and reduce reliance on human labor. In environments with agents, such as AGVs or pallets on conveyors, efficient collision-free path planning is essential. These path-planning problems for multiple agents are formalized as multi-agent path finding (MAPF) problems [1]. MAPF has been applied to various settings, including automated warehouses [2], airport taxiway management [3], automated parking systems [4], and video games [5].
In typical MAPF settings, agents are often sparsely distributed along aisles; however, increasing the agent density can improve space efficiency. For instance, high-density automated warehouses store significantly more goods than lower-density ones, reducing required floor space and operational costs (Fig. 1). Similarly, high-density automated valet parking systems allow vehicles to be parked and retrieved in a smaller area by minimizing the gaps between vehicles.
Despite these space-saving benefits, planning efficient paths using conventional techniques remains challenging. Typically, in high-density environments, only a subset of agents has active goals at any given time (e.g., moving some items for shipment). The remaining agents, although not assigned to any particular destination, still occupy space and obstruct the movement of the target agents. For the target agents to reach their objectives, the obstructing agents must be relocated efficiently. One approach is to use integer linear programming (ILP) to plan all agent paths jointly [6, 7]. However, this approach generally incurs significant computation time because all paths are optimized simultaneously by ILP. Another approach is to set temporary goals for obstructing agents and then apply existing MAPF methods. However, this often extends the paths of the target agents because the relocation of the obstructing agents is not optimized. Table I highlights these limitations and the need for a tailored solution to accommodate obstructing agents efficiently.
Figure 1: MAPF in a high-density warehouse environment. Only a subset of target agents (those with matching-colored flags) are assigned goals, whereas the rest act as movable obstacles. In addition to planning paths for target agents to reach their goals, efficiently relocating obstructing agents is crucial for successful navigation under high-density conditions.
@@
Methods agents agents efficiency
@@
This study refines MAPF for high-density environments (MAPF-HD) and develops an efficient algorithm tailored to this context. This algorithm employs phased heuristic techniques to relocate agents by swapping their positions with empty vertices (null agents). This approach, termed phased null-agent swapping (PHANS), solves MAPF-HD within a few seconds even in a densely packed environment, as shown in Fig. 2. The implementation is available at https://github.com/ToyotaCRDL/MAPF-in-High-Density-Envs.
@@ II-A Optimal Solvers
MAPF involves determining optimal collision-free paths for multiple agents [1, 8]. Conflict-based search (CBS) [9], a widely used optimal solver, plans agents’ paths to their goals (low-level search) and resolves collisions by branching (high-level search). However, CBS faces scalability challenges because MAPF is NP-hard [10]. Researchers have introduced improvements, such as Improved CBS, to prune the search more effectively [11]. Other optimal frameworks recast MAPF as a combinatorial optimization problem. For example, Lam et al. [12] applied the branch-and-cut-and-price ILP method. Similarly, Gómez et al. [13] used answer-set programming.
@@ II-B Suboptimal Solvers
Several bounded-suboptimal and heuristic algorithms have been developed to improve the scalability of MAPF [14]. Enhanced CBS (ECBS) [15] is a bounded-suboptimal solver that employs focal search. To enhance scalability, explicit estimation CBS (EECBS) [16] replaces the focal search in ECBS with explicit estimation search. Priority inheritance with backtracking (PIBT) [17], a decentralized MAPF algorithm, assigns dynamic movement priorities to agents and allows backtracking to resolve conflicts. Unlike static priority planning, PIBT enables agents to inherit and adjust priorities dynamically, preventing deadlocks and improving coordination in dense environments. Backtracking allows agents to reconsider past actions when conflicts arise, thereby ensuring flexible and adaptive navigation.
@@ II-C MAPF in High-Density Environments with Obstructing Agents
Even in dense environments, relocation problems can be solved when all agents have assigned destinations. However, to move target agents efficiently, the relocation of obstructing agents must be optimized.
Okoso et al. [6] introduced the cooperative automated valet parking problem (CoAVP). In CoAVP, agents are classified into three types: entering, departing, and relocating vehicles. Departing vehicles are assigned destinations. Entering and relocating vehicles can be anywhere in the parking lot, and their destinations are also optimized. Their ILP-based method extends the planar graph along the time axis and optimizes the paths within the time-expanded network. However, this approach presents two major challenges. First, ILP has high computational complexity and requires hundreds of seconds even in small environments (e.g., $15\times 15$ grids). Consequently, this method is impractical in real-world applications. Second, collision types specific to high-density environments have not been sufficiently investigated. To save parking space, agents must be densely positioned. In such environments, two adjacent agents must move in almost perfect synchrony, which may not be realistic. Following conflicts [1] should be considered for high-density environments.
@@
This study redefines MAPF in high-density environments and develops a heuristic approach to mitigate the high time complexity associated with ILP.
@@ III Multi-Agent Path Finding Problem in High-Density Environments
This section defines MAPF-HD. Unlike standard MAPF, which focuses on the paths of the target (goal-assigned) agents toward designated goals, MAPF-HD operates in environments where obstructing agents are densely packed. The objective of MAPF-HD is to minimize the number of time steps required for all target agents to reach their goals without collisions. Fig. 3 shows an example of MAPF-HD.
@@ III-A Environment
where $x$ and $x^{\prime}$ are row indices, and $y$ and $y^{\prime}$ are column indices. These indices correspond to the grid coordinates. The Manhattan distance $dst(v_{x,y},v_{x^{\prime},y^{\prime}})$ between two points is calculated as follows:
$dst(v_{x,y},v_{x^{\prime},y^{\prime}})=|x-x^{\prime}|+|y-y^{\prime}|.$ (3)
@@ III-B Agent
  • Target agents: $A^{\mathrm{tgt}}=\{a^{\mathrm{tgt}}_{1},\ldots,a^{\mathrm{tgt}}_{m}\}$, with $m>0$. Each target agent $a^{\mathrm{tgt}}_{i}$ is specified by a tuple $\langle s^{\mathrm{tgt}}_{i},g^{\mathrm{tgt}}_{i}\rangle$, where $s^{\mathrm{tgt}}_{i}\in V$ is the start vertex, and $g^{\mathrm{tgt}}_{i}\in V$ is the goal vertex.
  • Obstructing agents: $A^{\mathrm{obs}}=\{a^{\mathrm{obs}}_{1},\ldots,a^{\mathrm{obs}}_{n}\}$, with $n\geq 0$. Each obstructing agent $a^{\mathrm{obs}}_{i}$ is characterized solely by its start vertex, denoted as $\langle s^{\mathrm{obs}}_{i}\rangle$, as it does not have a goal.
@@ III-C Agent Action
Figure 3: Example of the MAPF-HD problem: (a) initial positions, (b) evacuation of obstructing agents, and (c) movement of the target agent.
Agents avoid collisions at each time step by managing the two conflicts:
@@
Some studies [9, 16, 17] are based on edge conflicts [1] rather than following conflicts. If following conflicts are not treated as conflicts, the agents must be spaced apart, and adjacent agents would need to move almost synchronously. However, in agent-dense environments, spacing the agents sufficiently apart may not be feasible. Moreover, perfect synchronization is not always realistic in real environments. Therefore, we treat following conflicts as prohibited. Note that following conflicts contain edge conflicts [1].
@@ III-D Solution and Objective
A solution to MAPF-HD is a set of collision-free paths $\{\pi_{1},\ldots,\pi_{m+n}\}$ ensuring that all target agents reach their goals at a certain time step $T$. None of the obstructing agents have goals, and they are freely positioned. Formally, the problem assigns a path $\pi_{i}=[\pi_{i}[0],\pi_{i}[1],\ldots,\pi_{i}[T]]$ to each agent, which must satisfy the following:
@@
The objective is to minimize makespan $T$, which is defined as the total number of time steps required for all target agents to reach their goals. In addition to makespan, classical MAPF often uses sum-of-costs [1], which is the sum of time steps required by each agent to reach its goal. In real-world applications such as warehouses and valet parking, the emphasis is on the speed of the target agents in reaching their goals, rather than on the total movement effort of all agents. Makespan captures this operational bottleneck, whereas sum-of-costs would be distorted by the large number of obstructing agents without goals.
@@ IV Proposed Method
Current methods [6, 7] use ILP to optimize the paths of all agents simultaneously, resulting in high computational complexity. To reduce the computational costs, we propose a two-stage heuristic approach termed PHANS. For the target agents, the first stage plans paths from their starting positions to their goals using the A* algorithm. The second stage sequentially evacuates the obstructing agents from these paths to clear the way for target agents. Section IV-A describes the methods used in both stages through an example with a single target agent, followed by a generalized algorithm for multiple target agents in Section IV-B.
@@ IV-A1 First Stage
To satisfy the second condition, the A* algorithm is modified by introducing an additional heuristic value, $h_{\mathrm{add}}$. This modification incorporates the expected waiting cost for obstructing agents to clear the path. The cost function $f$ is computed as the sum of $g,h$, and $h_{\mathrm{add}}$ as follows:
@@
$h_{\mathrm{add}}=\max(0,1+h_{\mathrm{evac}}-g).$ (7)
Here, $h_{\mathrm{evac}}$ estimates the cost of evacuating an obstructing agent by calculating its distance to the nearest empty vertex using (3). If the estimate is accurate and the obstructing agents are evacuated by the time the target agent arrives, $h_{\mathrm{add}}$ is 0. Otherwise, the waiting time is incorporated into $h_{\mathrm{add}}$. These computations are based on the initial positions; however, the set of empty vertices may change as obstructing agents are evacuated.
The proposed heuristic $h+h_{\mathrm{add}}$ is neither admissible nor consistent in general. When $h_{\mathrm{add}}=0$, it reduces to the standard admissible distance estimate. In other cases, it may overestimate the true evacuation cost, and its value can change nonmonotonically between neighboring states. Thus, the heuristic serves as a practical cost estimator that reflects the additional effort of relocating obstructing agents in high‑density environments, rather than a guarantee of optimality.
@@ IV-A2 Second Stage
    Plan a path from the null vertex to the obstructing agent (null-agent path), ignoring potential collisions with the other obstructing agents.
@@ IV-B Multiple Target Agents
For multiple target agents, the first stage can be executed simultaneously for all targets. Additionally, the second stage allows the simultaneous evacuation of multiple obstructing agents to enhance efficiency. The pseudocode for this generalized approach is provided in Algorithm 1. We perform sequential A* path planning using prioritized planning (PP), giving higher priority to targets with longer start-goal distances computed via (3).
Lines 1–4 in Algorithm 1 compute the distance from the start to the goal for each target agent in $A^{\mathrm{tgt}}$ using (3) and then prioritize the planning of target paths with the longest distances. As resolving conflicts during simultaneous path planning increases computational costs, paths are planned sequentially using PP.
@@
Lines 12–16 assign a null agent to each obstructing agent and plan the path between them. The null-agent assignment proceeds in order from the head of $B$, stopping if no null agents remain. In general, the nearest null agent is assigned to each obstructing agent. However, because target agents are prioritized over obstructing agents, any null agents lying on the target agent’s path between the target agent and the obstructing agent are reserved for the target agent. Therefore, these reserved null agents are not assigned to obstructing agents. Lines 18–21 move the target agent after the obstructing agent is evacuated. If there are no collisions, lines 22–25 execute the sequential relocation of the obstructing agents along the paths planned with their assigned null agents.
@@ IV-C How Fast Is PHANS?
Conversely, the methods proposed by Okoso et al. [6] and Makino et al. [7] extend the graph $G=(V,E)$ along the time axis and solve the minimum-cost flow problem through edges for path planning. With a time-axis size of $T$, their worst-case time complexity is $\mathcal{O}\left(2^{T|E|}\right)$.
@@
20:     bagt.path == null-agent path from bagt.goal to bagt.pos